I don't know how to do this question.
Factor the monomial of the following.
8a2b,30x2y2, 45m3n2,84b2c3d.
The 2's and 3's are to the power, but I don't know how to display that on the computer. I don't know what monomial is, either.

Answer by checkley75(3666)

8A^2B=4*2*A*A*B THESE ARE ALL THE FACTORS.
30*X^2Y^2=2*3*5*X*X*Y*Y.
45M^3N^2=3*3^5*M*M*M*N*N
84B^2C^3D=2*2*3*7*B*B*C*C*C*D